A microwave oven is on sale for 20% off
the regular price. The sale price is
$250.00. What is the regular price of the
microwave oven? How do i solve a question like this?

Answer by htmentor(1343)

Let P = the regular price
20% off means that the sale price is 80% of the regular price.
So we can write
0.80*P = 250
Solving for P gives
P = 250/0.80 = 312.50
So the regular price is $312.50